Please help me solve this equation
5/7-3 = y+7/2y-6 +1
My solution:
LCD of y-3 y-3(5/y-3)=y-3(y+7/2y-6)
=5*y+7/4
= 12/4
= 3

Answer by checkley77(12844)

5/7-3 = y+7/2y-6 +1
y+7y/2=5/7-3+6-1
(2y+7y)/2=5/7+2
9y/2=(5+2*7)/7
9y/2=(5+14)/7
9y/2=19/7
9y*7=2*19
63y=38
y=38/63
proof:
5/7-3=38/63+(7*38/63)/2-6+1
(5-3*7)/7=38/63+(266/63)/2-5
(5-21)/7=38/63+266/63*1/2-5
-16/7=38/63+266/126-5
-16/7=(38*2+266-126*5)/126
-16/7=(76+266-630)/126
-16/7=-288/126
-2016=-2016